Suppose the prices of the commodities A and B are p1 and p2 respectively. The consumer purchases X1 unit of A and X2 units of B. Suppose, the consumer has an income denoted by M and the consumer would spend the entire amount of M on these two commodities. How are p1, p2, X1, X2, and M related?


1
Expert's answer
2021-08-05T10:33:01-0400

Given prices, p1 and p2, find the quantities x1 and x2 which maximise consumer purchases

Necessary condition: 

"MRS=\frac{M\times U1}{M\times U2}=\frac{p1}{p2}\\MU1=1\\MU2=\frac{1}{x2}\\"

so

"MRS=x2"

Therefore the optimum occurs when

"x2=\frac{p1}{p2}"

Money left to buy x1 for:  

"m-p2\times x2=m-p2\times \frac{p1}{p2}=m-p1\\x1=\frac{m-p1}{p1}=\frac{m}{p1}-1"

if m>p1

b) Yes, if m>p1 he won’t buy any more beer when m increases.